OpenStudy (anonymous):

/ = strait line What values of a and b are a solution to the inequality /5-2a/ - b is less than 4? A. a = 6, b = -2 B. a = -4, b = 3 C. a = 3, b = -1 D. a = -3, b = 5

OpenStudy (anonymous):

\[/5-2a/ - b \le 4\]

OpenStudy (anonymous):

PLEASE HELP!!!

OpenStudy (anonymous):

Because there are two variables, the easiest way to solve this is to plug in the a and b values and check which one works.

OpenStudy (anonymous):

@trainwrecking I have not worked with the straight line things before.

OpenStudy (anonymous):

Oh, okay. The "straight lines" are actually the absolute value signs. They work a little like parenthesis, except after you solve what's inside them, it turns positive. For example, if a=6 and b=-2: \[\left| 5-2(6) \right| - (-2) \le? 4\]\[\left| 5-12 \right| +2 \le?4\]\[\left| -7 \right| +2 \le?4\] At this point you have |-7|. To get -7 out of the absolute value sign, you just have to make it positive. So now you'll have:\[7+ 2\le?4\] The answer ends up being that 9 is not less than or equal to 4, so A is NOT the correct answer choice. Do you think you can try the rest now?

OpenStudy (anonymous):

Ok thanks @trainwrecking I think I got it! :)

OpenStudy (anonymous):

Great! Just let me know if you need any more help :)

OpenStudy (anonymous):

OK @trainwrecking

OpenStudy (anonymous):

When doing inequalities with absolute values, it typically means you have to split things up. For example, if: \[|x-2| < y\]Becomes:\[-y< x-2<y \]

OpenStudy (anonymous):

ok @heril

OpenStudy (anonymous):

thanks